### v2

# AL-Bench: A Benchmark for Automatic Logging

Abstract.
Logging, the practice of inserting log statements into source code, is critical for improving software reliability. Recently, language model-based techniques have been developed to automate log statement generation based on input code. These tools show promising results in their own evaluation. However, current evaluation practices in log statement generation face significant challenges. The lack of a unified, large-scale dataset forces studies to rely on ad-hoc data, hindering consistency and reproducibility. Additionally, assessments based solely on metrics like code similarity fail to reflect real-world effectiveness. These limitations underscore the need for a comprehensive public benchmark to standardize evaluation.
This paper introduces AL-Bench, a comprehensive benchmark designed specifically for automatic logging tools. AL-Bench includes a high-quality, diverse dataset collected from 10 widely recognized projects with varying logging requirements and introduces a novel dynamic evaluation approach. Different from the existing evaluations that focus only on components of log statements like code similarity, AL-Bench assesses both the compilability of the code with inserted log statements and the effectiveness of the logs generated by them during runtime, which we believe can better reflect the effectiveness of logging techniques in practice. AL-Bench reveals significant limitations in the state-of-the-art tools. The codes with log statements generated by the state-of-the-art tools fail to compile in 20.1%-83.6% cases. In addition, even the best-performing tool only achieves 0.213 cosine similarity between the runtime logs produced by the generated log statements and the ground-truth log statements. The results reveal substantial opportunities to further enhance the development of automatic logging tools.
Logging, Software Maintenance, Logging Benchmark
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXccs: Software and its engineering
1. introduction
As software grows in size and complexity, logging has become increasingly essential to ensuring software reliability (He et al., 2021; Chen and Jiang, 2021). Logging means writing log statements into the source code, which generate runtime logs that record valuable information for a range of downstream tasks such as anomaly detection (Zhang et al., 2019; Lou et al., 2010; He et al., 2016; Du et al., 2017; Zhang et al., 2019), fault diagnosis (Zou et al., 2016), root cause analysis (Amar and Rigby, 2019; He et al., 2018b; Lin et al., 2016), and program verification (Ding et al., 2015; Shang et al., 2013). The effectiveness of these downstream tasks heavily relies on the quality of the software logs (He et al., 2018a). Therefore, appropriate logging is essential to capture critical behaviors during software operation (Yuan et al., [n. d.]).
Figure 1. An example of logging statement generation. Logging statement generation can be separated as three parts: determining the position, selecting the level, and specifying the message.
Numerous logging tools have been developed to assist software developers by automatically suggesting log statements based on provided code snippets (Mastropaolo et al., 2022; Liu et al., 2021; Xu et al., 2024; Xie et al., 2023). As illustrated in Figure 1, automatic logging typically includes three steps: (1) determining the position, (2) deciding the verbosity level, and (3) specifying the message to be recorded. Leveraging the advanced text generation capabilities of large language models (LLMs) (Floridi and Chiriatti, 2020; Raffel et al., 2020), LANCE (Mastropaolo et al., 2022) is the first end-to-end tool to integrate positioning, level selection, and message generation. UniLog (Xu et al., 2024) employed a warm-up and in-context learning (ICL) strategy to enhance performance. FastLog (Xie et al., 2023) improved the generation efficiency while maintaining precision. LEONID (Mastropaolo et al., 2023), based on LANCE, combined with deep learning and information retrieval technologies to enhance performance. SCLogger (Li et al., 2024) adapted static analysis to extend the context for the code snippet. These studies typically evaluate performance using ad-hoc data splitting from the entire dataset, focusing on metrics such as the accuracy of log statement components, including position, level, and message. Machine translation metrics, such as BLEU (Papineni et al., 2002) and ROUGE (Lin, 2004), are also employed to evaluate the quality of generated log messages. Although these evaluations offer valuable insights into logging tool performance, the use of low-quality data and incomplete assessments undermines the reliability of the results.
First, loose standards collection and inappropriate clean strategy compromise the quality of evaluation data and effectiveness in assessing the performance. Previous evaluation datasets were typically created by splitting the entire dataset. The data selection rules of the entire dataset are commonly loose standards to ensure sufficient data for training. These criteria fail to ensure the quality and consistency of the data. Moreover, to accommodate the limitations of the tools, they filtered out all instances exceeding 512 tokens, ignoring the long code snippets that are commonly encountered in real-world development environments. This approach undermines the overall effectiveness and real-world applicability of the evaluation results, as it does not accurately reflect the true complexity of software projects.
Second, the current evaluation method does not verify whether the generated log statements are compilable. Generating the compilable log statements is the first requirement when applying automatic logging tools in practice. To relieve developers from the heavy effort required to design and maintain log statements (Chen and Jiang, 2019, 2017), the basic requirement is to ensure that our tool can be seamlessly integrated into the DevOps process without introducing additional errors that require extra debugging effort from developers. Current evaluation methods merely focus on whether each component of log statements (i.e., position, verbosity level, message) matches the ground truth but cannot assess whether the generated log statements might introduce compilation errors such as wrong code format or use undefined variables. Evaluating the compilability of predicted log statements reflects the effectiveness and reliability of logging tools, which are essential for practical use.
Third, the evaluation method cannot evaluate the quality of runtime logs generated by the predicted log statements. Current evaluation methods assess the performance of tools based on the correctness of individual components of log statements. However, current metrics struggle to accurately reflect the quality of runtime logs in the real execution environment. Even a slight shift of log statement can lead to the miss of essential runtime details, such as a several-line shift from the ground truth or a mismatch in verbosity level. For instance, a minor difference in verbosity levels (e.g., debug vs. info) can cause critical information missed due to log level threshold settings in the source code. Therefore, we need to evaluate the quality of log statements in a real execution environment, with the goal of obtaining appropriate logs in specific scenarios, rather than merely focusing on the correctness of individual components or relying on statistical metrics to reflect the performance of logging tools.
Our Work. To address these challenges, we introduce AL-Bench: a comprehensive benchmark featuring a large-scale, diverse dataset and a novel approach for evaluating both the compilability and runtime logs produced by generated log statements. Our dataset comprises 42,224 instances collected from 10 popular, high-quality GitHub projects (GitHub, [n. d.]) spanning different domains with varying logging requirements, providing a robust foundation for evaluating automatic logging methods. Beyond the dataset, AL-Bench introduces a novel evaluation method called dynamic evaluation, which involves reintegrating generated log statements into real project code, followed by recompiling and executing them. This process allows for a realistic assessment of both compilability and runtime logs, highlighting major limitations in state-of-the-art tools: even the best tools fail to compile in 20.1% of cases, and the logs produced by generated log statements show only 0.213 cosine similarity to ground-truth logs. Through its rigorous and standardized evaluation approach, AL-Bench bridges the gap between real-world logging requirements and prior assessments, highlighting substantial opportunities to further advance the development of automatic logging tools.
This paper’s contributions are summarized as follows:
  1. (1)
    We collected a high-quality, diverse, and large-scale dataset comprising 42,224 instances from 10 popular, high-quality GitHub (GitHub, [n. d.]) projects, spanning various domains with differing logging requirements.
  2. (2)
    We propose a novel dynamic evaluation approach that assesses generated log statements in real-world settings by reintegrating them into projects to evaluate their compilability and resulting runtime logs.
  3. (3)
    We conducted a comprehensive evaluation of popular automatic logging tools and revealed the key limitations based on the analysis of the evaluation results.
  4. (4)
    All the data and code for AL-Bench are publicly available  [^1] https://github.com/shuaijiumei/logging-benchmark-scripts, providing valuable resources for both developers and researchers to advance the field of automatic logging.
2. background and motivation
This section introduces the overview of the current research on automatic logging tools, followed by the shortcomings in the evaluation work, and explains the motivation for AL-Bench.
2.1. Log Statement Generation
Logging, the process of generating informative log messages with appropriate verbosity levels at strategically placed locations within code, has long been recognized as a critical challenge in software engineering (He et al., 2021; Chen and Jiang, 2019, 2021; Chen et al., 2021). Over the years, substantial research efforts have aimed to support developers in crafting more effective logging statements, which in turn enhance software maintenance and testing (Jia et al., 2018; Zhao et al., 2017; Zhu et al., 2015). Early studies in this domain often addressed isolated subproblems, typically operating under stringent assumptions that limit the applicability of their findings in real-world scenarios. For example, Li et al. (Li et al., 2021) proposed DeepLV to predict the appropriate logging level by taking surrounding code features into a neural network. Liu et al. (Liu et al., 2022) proposed Tell to further adapted flow graphs to help the suggestions of verbosity levels. Zhu et al. (Zhu et al., 2015) proposed LogAdvisor and Yao et al. (Yao et al., 2019) proposed Log4Perf to assist developers to add new log statements in a specific position. Ding et al. proposed LoGenText (Ding et al., 2022) and LoGenText-Plus (Ding et al., 2023) to advise developers what should be logged, and Liu et al. (Liu et al., 2021) proposed tools for deciding which variables should be logged. However, none of them can generate a complete log statement.
Recently, Mastropaolo et al. (Mastropaolo et al., 2022) proposed the first end-to-end tool LANCE to generate complete log statements based on T5 (Raffel et al., 2019). Following LANCE, Xu et al. (Xu et al., 2024) proposed UniLog to adapt ICL and warm-up strategy to enhance the LLM ability for generating log statements. Xie et al. (Xie et al., 2023) proposed FastLog increase the generation time while keeping the accuracy, and Mastropaolo et al. (Mastropaolo et al., 2023) further proposed LEONID with a combination of Deep Learning (DL) and Information Retrieval (IR) achieving a better performance.
While end-to-end automatic logging tools have demonstrated promising results in their respective evaluations, our analysis reveals notable issues in both the datasets used and the evaluation methods employed.
Figure 2. The low-quality examples in the LANCE dataset follow the strictest data collection rules.
2.2. Limitations of Evaluation Methodology
2.2.1. Evaluation Data
Current evaluation datasets are ad-hoc, derived from splitting the entire collected data according to loose standards. More specifically, to ensure a sufficient amount of training data, previous studies had to adopt lenient rules for data collection. Additionally, some tools, to accommodate the limitations of their backbone models, further filter out data exceeding 512 tokens in length. These strategies compromise data quality, reducing the reliability and effectiveness of evaluation results. For example, LANCE adopted the strictest data collection rules among all tools, requiring a minimum of 500 commits, 10 contributors, and 10 stars, while also excluding forked repositories on GitHub (GitHub, [n. d.]). However, even in the LANCE dataset, low-quality data is pervasive. As shown in Figure 2, several patterns of low-quality logging practices persist. These patterns include the duplication of variables within a single log statement, creating unnecessary redundancy; the inclusion of empty strings or meaningless content, resulting in uninformative messages; excessive use of special characters or punctuation, making printed logs difficult to parse; and mismatched logging levels, where critical messages are assigned inappropriate severities, leading to misclassification (Adzic, 2006; Chen and Jiang, 2017). Using low-quality data as an evaluation dataset fails to produce reliable results. Such results do not align with the expectation that logging tools should generate high-quality log statements within the code. Furthermore, to accommodate the input length limitations of these tools, previous studies (Mastropaolo et al., 2022, 2023; Xu et al., 2024; Xie et al., 2023) have filtered out instances longer than 512 tokens. While this data-cleaning strategy reduces the complexity of evaluation data, it also introduces significant biases by excluding longer code snippets, which are common in real-world development environments. This approach simplifies the evaluation process but fails to capture the full scope of challenges that logging tools would face when dealing with complex and extended code bases. Consequently, the evaluation results may not accurately reflect the tools’ ability to handle the demands of real-world software development, where longer snippets and more intricate code structures are prevalent. Therefore, there is a pressing need for a public, large-scale, high-quality, and diverse benchmark dataset that can better represent real-world codebases and provide a standardized platform for evaluating automatic logging tools.
Figure 3. An example of the limitation of static evaluation, where only the quality of log statements is considered. It cannot assess the compilability of the code or quality of the printed runtime logs.
2.2.2. Evaluation Method
The current evaluation method focuses on assessing model performance by comparing the accuracy of each log statement component (i.e., position, verbosity level, message). We refer to this as static evaluation in this paper. Figure 3 provides an example of static evaluation to demonstrate how it works. In this example, only the verbosity level matches, so the matched level count will increase by one, and BLEU scores will be used to calculate the average score. Finally, the matched portion of each component and the average BLEU score will be combined to reflect the performance of the tools. These statistical metrics provide insights into evaluating the quality of predicted log statements; however, this approach still does not align with the goal of ensuring that the predicted log statements generate high-quality logs.
Figure 4. An example of compilation failure: Unreachable Statements
First, it cannot assess the compilability of predicted log statements. Compilability is a fundamental requirement for generating appropriate logs. This capability ensures that log statements can be applied to the source code without introducing compilation errors, avoiding wasting developers’ efforts to debug for the predicted log statements. However, the current evaluation methodology, which focuses solely on comparing the individual components of log statements, cannot identify predicted log statements that use undefined variables or contain incorrect code syntax which could introduce compilation errors. As shown in Figure 4, the predicted log statement is injected into an unreachable position. Applying static evaluation would fail to recognize that this prediction could cause compilation errors. Instead of classifying it as a completely incorrect prediction, the evaluation would count it as a level match instance and assign it a BLEU score of 55.03, contributing to the statistical value. This could introduce significant bias when evaluating the performance, appearing favorable in metrics but performing poorly in actual usage.
Second, it cannot assess the runtime logs generated by predicted log statements. The goal of writing log statements is to obtain the appropriate logs during software execution so that the quality of log statement is ultimately decided by the quality of runtime logs. Although comparing the correctness of log statement components provides insights into assessing the quality of logs, it can introduce bias when using statistical metrics that focus solely on log statements to evaluate the performance of logging tools. A simple example is a mismatched verbosity level between the predicted log statement and the ground truth. In a case where only the verbosity level is mismatched while the other components are exactly matched, this instance would receive a high score in static evaluation. However, in a real-world scenario, for important error logs, if the predicted level is below the threshold, the logs will not be generated, regardless of how critical the message is. Conversely, for debug logs, if the predicted level exceeds the threshold, these logs could be printed in the production environment, leading to issues such as sensitive information leakage and increased storage costs. These two situations demonstrate that even a shift in verbosity level can turn a log statement into a completely bad case. A more complicated case is presented in Figure 5. Under static evaluation, this prediction with only a shift in two lines in position, also could be rewarded highly in statistic metrics. However, it is completely incorrect: placing the log statement at line 2 risks a NullPointerException if the timer object is null when timer.get() is called. This would cause the program to crash during execution, meaning that not only would the logs fail to be generated, but the entire program could terminate unexpectedly. In real-world scenarios, this kind of error is critical and goes unnoticed in static evaluations. While static evaluation rewards this prediction for structural correctness, it fails to account for the actual behavior of the code when executed. Therefore, this example reinforces the need for more comprehensive evaluation methods that go beyond code-level metrics and consider the real-world implications of log statements.
In conclusion, a log statement’s quality should be determined by its ability to avoid introducing extra errors and generate meaningful, contextually appropriate logs.
Figure 5. An example of the limitation of static evaluation. The prediction differs from the ground truth by a slight upward shift of two lines. It causes a NullPointerException when timer is null and fails to generate logs.
Insights: The motivating study underscores major limitations in current evaluation efforts, including the lack of a diverse, high-quality dataset and methods that do not meet real-world needs. To improve log statement evaluation, we should collect a high-quality, diverse dataset and use an execution-based method to directly evaluate the runtime logs generated by log statements.
3. AL-Bench
In this section, we introduce our benchmark, AL-Bench, which builds on insights from previous studies (Mastropaolo et al., 2022, 2023; Xu et al., 2024; Xie et al., 2023). AL-Bench consists of two key parts: static evaluation and dynamic evaluation, both of which we will introduce in detail. The static evaluation component includes a high-quality dataset with 42,224 instances and five metrics, based on the static evaluation method outlined in Section 2.2.2. The dynamic evaluation component introduces a separate dataset with 2,238 instances and four metrics, following the dynamic evaluation approach, which will be described later in this section.
3.1. Static Evaluation
3.1.1. Static Evaluation Method.
Static evaluation focuses on log statement components—position, verbosity, and message—and has been the primary method in prior studies (Mastropaolo et al., 2022, 2023; Xu et al., 2024; Xie et al., 2023), with details provided in Section 2.2.2.
3.1.2. Dataset Construction
As discussed in Section 2.2.1, the quality of the evaluation dataset is crucial for assessing the performance of tools. The strictest existing rules for dataset collection—500 commits, 10 contributors, and 10 stars—are inadequate to ensure the quality of the evaluation dataset. Drawing inspiration from the use of GitHub repository stars as an effective metric for identifying high-quality code datasets  (Jiang et al., 2024), and considering the unique characteristics of log statements, we propose three critical criteria for dataset selection: repositories must have at least 10,000 stars, 1,000 log statements, and 500 log-related issues to qualify for inclusion in our dataset. These standards guarantee well-maintained, widely used projects where developers prioritize high-quality log statements. Given the diverse requirements for log statements in different scenarios, we ultimately selected 10 projects based on domain diversity, industrial applications, and their relevance to prior logging studies (Chen and Jiang, 2017, 2019; Li et al., 2018, 2019; Ding et al., 2022, 2023; Li et al., 2020). As shown in Table 1, our final dataset includes projects with a total of 22,787 code snippets and 42,224 log statements, covering a wide range of logging needs and practices. The dataset spans multiple domains, including database management, task scheduling, distributed storage, messaging systems, and IoT platforms, ensuring diversity in logging scenarios. Each domain presents unique requirements for log statements.
For example, database management systems such as DBeaver and Doris prioritize minimizing the impact of logging on high performance. Task scheduling systems, including DolphinScheduler, rely on logging to trace task dependencies and monitor runtime statistics. Similarly, distributed systems like Hadoop and Zookeeper require robust logging practices to address challenges in distributed coordination, fault tolerance, and scalability, which differ from the requirements of other command logs. Messaging systems such as Kafka and Pulsar have adapted logging practices to trace message flows, ensure reliable message delivery, and debug asynchronous communication. Meanwhile, IoT platforms like ThingsBoard utilize logging to manage device connectivity, monitor data streams, and enable real-time system oversight. Identity and access management systems such as Keycloak prioritize protecting sensitive information in logs to prevent privacy breaches. By including diverse projects across these domains, the dataset ensures comprehensive coverage of different logging practices, making it suitable for benchmarking and analyzing log-related tasks under various operational contexts.
In addition to emphasizing data quality, we also addressed the potential risk of data contamination. Since all our data were extracted from public GitHub repositories, which may have been used for training pre-trained models, we implemented precautions to minimize this risk. Specifically, we collected the latest version of each project to reduce the likelihood of it being included in any model’s pre-training data. Furthermore, we wrapped the code snippets in a common class and standardized the formatting using Google-Java-Format (Google, [n. d.]). This approach altered the format of the code to prevent pre-trained models from recognizing the same information and structure. These strategies have been demonstrated as effective in recent studies (Xu et al., 2024; Tian et al., 2024). Although we adopted effective strategies, data contamination cannot be entirely avoided (Cao et al., 2024). However, our methods minimize this risk and have been proven effective in previous work, providing a solid foundation for analysis and evaluation. In the future, we plan to regularly update our dataset to ensure that the evaluation data remains current. After completing the necessary appeal actions, we finalized our static evaluation dataset.
Table 1. Details of AL-Bench dataset.
Dataset Domain Code snippets with log statements Total log statements
Dbeaver Database Management 1,182 1,725
Dolphinscheduler Task Scheduling 898 1,918
Doris High Performance Database 1,686 2,965
Flink Data Processing 1,956 3,268
Hadoop Distributed Storage 8,751 16,002
Kafka Messaging Systems 1,706 3,421
Keycloak Identity and Access Management 533 1,004
Pulsar Messaging Systems 3,590 7,254
Thingsboard IoT Platform 1,601 2,731
Zookeeper Distributed Coordination 884 1,936
Total - 22,787 42,224
3.1.3. Metrics
Building on previous studies (Mastropaolo et al., 2022, 2023; Xu et al., 2024; Xie et al., 2023), We adopted five metrics for static evaluation. In addition to the previously established metrics, we propose two new ones: Dynamic Variable Accuracy and Static Text BLEU, designed to better reflect the real-world quality requirements of log statements.
Metric 1: Level Accuracy (LA): Level Accuracy focuses on log level values, which indicate the importance of the message and are a crucial part of log statements. Some common levels, like ”info,” refer to the normal information of runtime behavior. ”Warning” indicates potential problems that might not immediately cause disruption but could lead to future issues if not resolved. ”Error” refers to runtime anomalies or issues that need to be addressed. Each level refers to different meanings so for Level Accuracy we strictly compared the level of predictions and source code and used the exactly matched level number $L_{c}$ divided by the total number of log statements $N_{a}$ to get the value of this metric: $LA=\frac{L_{c}}{N_{a}}$.
Metric 2: Position Accuracy (PA): Position Accuracy focuses on the precise location of log statements within the source code, which is crucial for tracking and debugging software behavior. Correct placement of log statements helps in accurately tracing the execution flow and diagnosing issues. For Position Accuracy, we rigorously compare the predicted positions of log statements with their actual positions in the source code. This metric is calculated by taking the number of correctly positioned log statements $P_{c}$ and dividing it by the total number of log statements $N_{a}$ to obtain the accuracy value: $PA=\frac{P_{c}}{N_{a}}$.
Metric 3: Message Accuracy (MA): Message Accuracy evaluates how accurately the predicted log messages match the ground truth log messages, which is essential for providing meaningful and relevant information during runtime. The content of log messages helps developers understand the system’s behavior, and inaccuracies in message generation can lead to confusion or missed insights during debugging. For Message Accuracy, we compare the predicted log messages to the actual messages in the source code. This metric is calculated by determining the number of log messages that are fully identical to the ground truth $M_{c}$ and dividing it by the total number of log messages $N_{a}$, yielding the accuracy value: $MA=\frac{M_{c}}{N_{a}}$.
Metric 4: Dynamic Variable Accuracy (DVA): Dynamic Variable Accuracy focuses on the dynamic variable in log message, for example, (”The server is running, {}”, status), in this message, status is regarded as the dynamic variable in log message which might vary according to the runtime behaviors of software. And in the log message (”The server run on the ports, {}”, args.status ? localPort : remotePort), the whole expression (args.status ? localPort : remotePort) is reckoned as the dynamic variable which could decide the port information recorded in log files. We aim to use this metric to ensure that the dynamic information recorded in logs remains consistent. We extract the dynamic variables from both the source code and predictions and then compare them to record instances where they match exactly. This metric is calculated by taking the number of the exactly matched $DP_{c}$ and dividing it by the total number of log statements $N_{a}$ to obtain the accuracy value: $DVA=\frac{DP_{C}}{N_{a}}$.
Metric 5: Static Text BLEU (STB): Static Text BLEU focuses on the static part of the log message. Unlike the dynamic variable, the static part always records the same information in log files, which will not vary due to the runtime behavior of software. For example, in the log message (”The server is running, {}”, status), ”The server is running, {}” is regarded as the static part. Since this part primarily consists of natural language content, we use BLEU (Papineni et al., 2002) metric to evaluate the quality of the static text. In our implementation, we use the BLUE-DM variant (Chen and Cherry, 2014; Shi et al., 2021), i.e., the sentence-level BLEU without any smoothing method.
3.2. Dynamic Evaluation
Figure 6. The general workflow of dynamic evaluation. First, compile the project and run the unit test to obtain ground truth logs. Then, replace log statements with predictions, rerun the test to generate new logs, and finally analyze the results.
3.2.1. Dynamic Evaluation Method
Different from static evaluation, dynamic evaluation focuses on compiling the code and runtime-generated logs, addressing static evaluation’s inability to verify code compilability and runtime logs. To directly assess runtime logs, we generate them using unit tests, which are widely used in software development to verify code functionality in isolated scenarios. Unit tests are readily available in most projects and they are designed to test the functionality and behaviors of code when facing different situations. They offer a natural method for simulating realistic situations, allowing generating logs without the need for complex runtime environments. Figure 6 demonstrates the general workflow of dynamic evaluation. The process begins with compiling the source code and executing the unit tests to obtain logs from the original log statements. Using tools like Jacoco (Jacoco, [n. d.]) and SureFire (surefire, [n. d.]), we collect the logs and remove the log statements covered by unit tests in the source code. Next, we input the modified source code into an automatic logging tool to generate predicted log statements. Then we inserted the predicted log statements back into the source code, replacing the original log statements. After recompiling the modified code, we rerun the unit tests to capture the logs produced by the inserted predicted log statements. The whole process provides us with two sets of logs—those generated by the original log statements and those generated by the predicted log statements. Finally, we evaluate the effectiveness of the predicted log statements using two key metrics: Compilation Success Rate and Log File Similarity. Compilation Success Rate ensures that the predicted log statements do not introduce compilation errors, while Log File Similarity measures the similarity between the logs generated by the predicted log statements and those generated by the original log statements. In the following sections, we will detail how we built the dataset for dynamic evaluation and introduce the specific metrics used to measure the performance of the automatic logging tools.
3.2.2. Dataset Construction
To build the dynamic evaluation dataset, we begin by compiling the entire project to ensure all dependencies are resolved and the project is ready for execution. Next, we systematically identify all available unit tests within the project. For each unit test, we execute it individually while employing the Jacoco Plugin (Jacoco, [n. d.]) to trace code coverage, specifically identifying whether the unit test interacts with or covers any log statements in the codebase. Simultaneously, we use the SureFire Plugin (surefire, [n. d.]) to capture the logs generated during the execution of the unit tests.
By correlating Jacoco coverage data with SureFire logs, we can match specific code snippets containing log statements to the corresponding unit tests that cover them, along with the runtime logs they generate. This process enables us to construct a comprehensive dataset consisting of triples: the code snippet, the unit test that triggers it, and the recorded logs. These triples are critical for dynamic evaluation, as they provide the ground truth for assessing the quality and effectiveness of predicted log statements in actual runtime scenarios.
To construct a dynamic evaluation dataset, it is essential to select high-quality projects with comprehensive unit tests, as these tests provide a realistic simulation of diverse production environments. Each instance of dynamic evaluation requires recompiling the project and executing the test to collect logs, making the process highly time-consuming. To balance this intensive time requirement with the need for sufficient dataset diversity and quantity, we employed Hadoop as our dynamic evaluation platform. This approach allowed us to build a dataset of 2,238 instances that balances diversity with sufficient size. Additionally, we open-source the entire suite of tools used in this evaluation process, enabling researchers and organizations to easily deploy and customize their own dynamic evaluation datasets.
3.2.3. Metrics
In dynamic evaluation, we proposed four metrics to assess the performance of logging tools: Compilation Success Rate, Log Similarity, False Positive Log Generation Rate, and False Negative Log Generation Rate, which we will introduce below.
Metric 6: Compilation Success Rate (CSR): Compilation Success Rate measures the syntactic correctness of the predicted log statements. As discussed earlier, after replacing the original log statements with the predictions, we recompile the project. However, due to issues such as undefined variables in the predictions or missing/outdated dependencies in the project environment, not all predictions can be successfully compiled. We recorded the successfully compiled code snippet number as $C_{s}$ and all code snippets as $C_{a}$. The metric is then calculated as: $CSR=\frac{C_{s}}{C_{a}}$.
Metric 7: Log Similarity Metrics Group (Cosine Similarity, BLEU, ROUGE): This metric evaluates how closely logs generated by predicted statements match those produced by ground truth statements. To eliminate unnecessary differences, we remove log headers (e.g., timestamps), retaining only log content. For a comprehensive assessment, we apply multiple similarity measures, including Cosine Similarity (Salton et al., 1975), BLEU (Papineni et al., 2002), and ROUGE (Lin, 2004). Cosine Similarity, commonly used in text analysis, calculates the cosine of the angle between two TF-IDF (Jones, 2021) vectors, yielding 1 for identical vectors and 0 for orthogonal ones. Using TF-IDF, we down-weight frequent terms, emphasizing distinctive content in logs. This method effectively captures the similarity between meaningful log content, filtering out redundant information for a more accurate relevance measure. ROUGE, on the other hand, focuses on recall by comparing n-grams between the predicted and reference logs. It evaluates how much of the reference content is preserved in the prediction. The most commonly used variant is ROUGE-N, which calculates the overlap of n-grams between two texts.
Metric 8: False Positive Log Generation Rate (FPLR): This metric measures the proportion of predicted log statements that generate logs during unit test execution when the ground truth log statements would not have produced any logs. It helps assess whether the predicted log statements introduce unnecessary or redundant logs in scenarios where no log should be generated. The number of false positive instances is recorded as $FP$, and the total number of predictions is $P$. the metric is calculated as: $FRLR=\frac{FP}{P}$.
Metric 9: False Negative Log Generation Rate (FNLR): This metric evaluates the proportion of predicted log statements that fail to generate logs during unit test execution when the ground truth log statements should have produced logs. It highlights instances where the predicted logs miss important events or information. The number of false negative instances is recorded as $FN$, and the total number of predictions is $P$. The metric is calculated as: $FNLR=\frac{FN}{P}$.
4. EXPERIMENTS
In this section, we use AL-Bench to evaluate existing end-to-end automatic logging tools and analyze the evaluation results to find insights to guide the following work. From the perspective of empirical software engineering, we set three research questions:
  • RQ1: How well are logging tools at predicting log statements in static evaluation?
  • RQ2: How well can the predicted log statements be compiled successfully?
  • RQ3: How well do the generated log statements print logs in dynamic evaluation?
Specifically, RQ1 examines how well logging tools predict log statements using the AL-Bench static evaluation dataset and corresponding static evaluation method. This research question aims to evaluate the accuracy and effectiveness of the tools in generating log statements that closely resemble the ground truth. RQ2 focuses on assessing the compilability of the predicted log statements, determining whether these predictions can be seamlessly integrated into the code without causing compilation errors. Finally, RQ3 evaluates the runtime logs produced by the generated log statements. We examine not only the generation of inappropriate logs but also the similarity between the predicted logs and the expected logs in cases where logs are correctly generated. We begin by introducing the automatic logging tools selected for evaluation. We will provide a detailed analysis of these tools, focusing on their performance across the three research questions. Each tool is assessed for its ability to predict accurate log statements, compile them successfully, and generate meaningful logs. This comprehensive evaluation allows us to draw insights into the strengths and limitations of current approaches to automatic logging.
4.1. Automatic Logging Tools
Automatic logging is a hot topic, leading to the development of many tools for determining specific parts of log statements and end-to-end automatic logging tools in recent years. In this paper, we focus on end-to-end logging tools. We reached out to the authors of popular end-to-end automatic logging tools for assistance in rebuilding these tools. Because SCLogger (Li et al., 2024) is still under construction, we ultimately selected four methods for evaluation: LANCE (Mastropaolo et al., 2022), LEONID (Mastropaolo et al., 2023), FastLog (Xie et al., 2023), UniLog (Xu et al., 2024). We detailed the method in the following.
LANCE: $LANCE$ (Mastropaolo et al., 2022) is the first model designed to generate and insert complete log statements in code. It takes a method requiring a log statement and outputs a meaningful log message with an appropriate logging level in the correct position. Built on the Text-To-Text Transfer Transformer model, $LANCE$ is trained specifically for injecting proper logging statements.
LEONID: $LEONID$ (Mastropaolo et al., 2023) is the updated version of $LANCE$. With a combination of DL and Information Retrieval (IR), $LEONID$ achieved a better performance. $LEONID$ provided two versions, $LEONID_{S}$ is for single log statement generation, and $LEONID_{M}$ is for multiple log statements generation. Since $LEONID_{M}$ can generate more than one log statement at a time, it introduces ambiguity in determining the correct correspondence between the generated and expected log statements when more than one log statements are generated by static evaluation (Mastropaolo et al., 2023). Therefore, we only applied $LEONID_{S}$ for static evaluation.
UniLog: $UniLog$ (Xu et al., 2024) is the first attempt to adapt Warm-up and In-context-learning strategy to enhance the model’s ability to generate log statements. Due to limitations in assessing the original UniLog, we reproduced it using two backbone models: CodeLlama-7B (Rozière et al., 2023) and DeepSeek-V3 (DeepSeek-AI et al., 2024). We applied the warmup process exclusively to the CodeLlama backbone model, while employing the ICL strategy to construct prompts for both models. The data are sourced from LANCE (Mastropaolo et al., 2022) to warm up and generate ICL content. The effectiveness of In-Context Learning often depends on whether the examples are in-distribution or out-of-distribution relative to the evaluation data. Since LANCE’s data distribution differs from our evaluation data, this may affect UniLog’s performance. We will use $UniLog_{{cl}}$ to represent the version based on CodeLlama-7B, $UniLog_{{ds}}$ to represent the version based on DeepSeek-V3.
FastLog: FastLog (Xie et al., 2023) defines the logging task in two steps: finding the position and generating and inserting a complete log statement into the source code. This approach avoids rewriting the source code, a key limitation of $LANCE$. They utilized PLBART (Ahmad et al., 2021) as the base model to fine-tune two separate models: one for predicting insertion position, the other for generating log statements. With the heuristic rule, log statements only appear after certain special characters, FastLog enhances efficiency while maintaining accuracy in generating log statements.
Table 2. Static evaluation on the complete logging task was conducted. The values in the Original lines indicate the reported results from the respective methods’ own evaluations. The best performance evaluated by AL-Bench across different metrics is highlighted.
Method PA LA MA DVA STB (BLEU)
Now Original Now Original Now Original
FastLog 57.54 $58.84$ 62.72 $59.75$ 6.90 $4.52$ 17.66 20.20
$UniLog_{{ds}}$ $36.31$ $76.90$ $36.31$ $72.30$ $5.19$ $22.40$ 15.73 11.60
$UniLog_{{cl}}$ $23.29$ $76.90$ $23.29$ $72.30$ $2.31$ $22.40$ 16.08 8.43
LANCE $34.67$ $65.40$ $34.67$ $66.24$ $2.99$ $16.90$ 14.42 6.52
$LEONID_{S}$ $15.22$ $76.45$ $15.22$ $73.53$ $1.38$ $31.55$ 5.35 2.03
4.2. RQ1: How well are logging tools at predicting log statements in static evaluation?
Figure 7. Performance of logging tools among different projects. The performance of each tool varies considerably across different projects and the trends of all methods across all projects generally remain consistent.
We evaluated four automatic logging tools using the AL-Bench static evaluation dataset and corresponding static evaluation methods, ensuring a fair comparison of their performance.
Following previous studies (Mastropaolo et al., 2022; Xu et al., 2024; Xie et al., 2023; Mastropaolo et al., 2023), we first processed our dataset to create evaluation pairs $<M_{s},M_{t}>$ with $M_{s}$ representing the input provided to the model (i.e., $M_{s}$ with one removed log statement) and $M_{t}$ being the expected output (i.e., $M_{t}$ is the removed log statement). After processing, we got 42,224 instances. The evaluation results are presented in Table 2. The metrics for Position Accuracy (PA), Level Accuracy (LA), and Message Accuracy (MA) are adapted from the evaluation frameworks used by the selected logging tools. Therefore, we label the values of these metrics as reported in their respective evaluation studies. The results demonstrate significant performance differences among the four logging tools.
FastLog consistently outperforms the other tools across all metrics, but it still faces challenges in specific areas. For instance, it scores only 17.66 in Dynamic Variable Accuracy (DVA) and 20.44 in Static Text BLEU (STB), suggesting that even the best tool struggles to generate meaningful natural language descriptions and accurately select dynamic variables for logging. Despite these shortcomings, FastLog shows only minor deviations from previous studies, maintaining stable performance across all metrics. In contrast, the other three tools, particularly LEONID, experience a significant decline in accuracy across all evaluated metrics.
As illustrated in Figure 7, the performance of each tool varies considerably across different projects. This fluctuation suggests that the reliability and stability of logging tools remain inconsistent when applied to diverse scenarios. The performance fluctuations of logging tools across different projects are due to variations in project complexity and the diverse logging requirements. Additionally, limited training data diversity may hinder the tools’ ability to perform consistently across unfamiliar scenarios. These factors collectively contribute to the observed instability in logging tool performance across different projects. As demonstrated in our results, the variability in tool performance across different projects underscores the necessity of diverse evaluation data to capture the full spectrum of a tool’s capabilities and limitations.
To investigate the causes of performance differences compared to previous studies, we observed that earlier evaluation datasets were filtered to exclude instances longer than 512 tokens. To assess the impact of this filtering, we divided our dataset into two groups: instances with shorter than 512 tokens and those with longer than 512 tokens. This division enables us to determine whether instance length contributes to the performance discrepancies observed among the logging tools, providing further insight into how different data characteristics influence tool effectiveness.
Table 3. The performance of tools when facing different length input data. Long means data longer than 512 tokens, Short means data shorter than 512 tokens, and $\Delta$ means the difference in logging tools performance when facing the Short data and the Long data.
Method Dataset PA LA MA DVA STB
FastLog Long 43.36 64.92 7.02 11.98 20.97
Short 61.79 62.33 6.88 19.36 19.98
$\Delta$ $\downarrow 18.43$ $\uparrow 2.59$ $\uparrow 0.14$ $\downarrow 7.38$ $\uparrow 0.99$
$UniLog_{d}{}_{s}$ Long 22.28 66.66 8.82 15.56 17.05
Short 40.51 59.52 4.11 15.78 9.97
$\Delta$ $\downarrow 18.23$ $\uparrow 7.14$ $\uparrow 4.11$ $\downarrow 0.22$ $\uparrow 7.08$
$UniLog_{c}{}_{l}$ Long 8.70 51.62 2.84 12.91 8.20
Short 27.67 51.78 2.62 17.02 8.80
$\Delta$ $\downarrow 18.97$ $\downarrow 0.16$ $\uparrow 0.22$ $\downarrow 4.12$ $\downarrow 0.60$
LANCE Long 0 0 0 0 0
Short 45.05 47.61 3.89 18.74 8.12
$\Delta$
$LEONID_{S}$ Long 0 0 0 0 0
Short 19.78 21.46 1.79 6.95 2.64
$\Delta$
After dividing the dataset, we obtained two groups: a longer dataset containing 9,732 instances and a shorter dataset with 32,492 instances. The evaluation results in Table 3 show a clear difference in tool performance based on the length of the instances. Notably, LANCE and LEONID struggled with instances longer than 512 tokens, failing to generate syntax-correct code and, in some cases, producing incomplete code. This explains why their scores for these cases are reported as zero, highlighting the input length limitations of both tools and their inability to handle longer, more complex instances effectively. $UniLog_{cl}$ and FastLog show a considerable drop in PA and DVA when handling longer data, indicating that they struggle to predict log positions and select dynamic variables in more complex instances. $UniLog_{ds}$also shows a dramatic drop in position accuracy, but only a slight decrease in the ability to decide dynamic variables for recording. However, with more sufficient information and a more powerful backbone model, it demonstrates significantly better performance in other metrics. This indicates that with a more powerful backbone, UniLog can better understand the context with longer inputs. Although $UniLog_{d}s$ is equipped with a more powerful backbone, it still struggles to choose the appropriate position for the log statement. As widely demonstrated that the LLMs are not good at counting numbers (Ahn et al., 2024), we might change the output format by directly adding log statements into the source code rather than generating the exact line number. Furthermore, by simply analyzing the control flow graph of the code, we might be able to exclude positions where logging is not feasible to leave fewer choices.
Answer to RQ1. FastLog performs best in AL-Bench ’s static evaluation but still faces challenges in generating meaningful descriptions and selecting key dynamic variables. All tools struggle to maintain stable performance under varying logging requirements and complex data, especially in determining log positions. Since LLMs struggle with counting, it might be better to avoid outputting exact line numbers and instead add tags in the source code. Additionally, analyzing the control graph to exclude impossible positions might be useful to enhance log position determination.
Table 4. Compilation Failure Rates Across Methods.
Methods Compilation Success Rates
FastLog 79.9%
$UniLog_{{cl}}$ 70.3%
$UniLog_{{ds}}$ 60.2%
LANCE 49.4%
$LEONID_{M}$ 25.0%
$LEONID_{S}$ 16.4%
Table 5. Compilation Failure Reasons Analysis.
Failed Reason Failed Number
Using Wrong Logging Name 56
Using Undefined Method 21
Using Undefined Variables 15
Incompatible Types 3
Unreachable Statements 1
Other 4
Total 100
4.3. RQ2: How well can the predicted log statements be compiled successfully?
To evaluate the tools’ ability to generate compilable log statements, we replace existing logs with predicted ones in the dynamic evaluation dataset and recompile the project to check for successful compilation. As shown in Table 5, the best-performing tool, FastLog, achieves a 79.9% Compilation Success Rate, followed by $UniLog_{c}l$ at 70.3%, $UniLog_{d}s$ at 60.2%, LANCE at 49.4%, $LEONID_{M}$ and $LEONID_{S}$ at 25.0% and 16.4%, respectively. The results reveal a key limitation of LANCE and LEONID that they regenerate the entire code snippet, which increases the risk of unintended code changes and can potentially lead to compilation errors. In contrast, FastLog and UniLog focus solely on generating the new log statement, minimizing the risk of errors by limiting codebase modifications. Although FastLog is one of the best tools available, it still leads to significant instances of compile failure. When log statements fail to compile, they can prevent the system from functioning as intended and complicate debugging or error tracking. To ensure the reliability of automatic logging tools in real-world environments, it’s crucial that the log statements they generate integrate smoothly into the existing codebase.
Given the importance of reliable compilation, we conducted a manual review of the compilation failures to identify the specific causes. This analysis will provide further insight into the key weaknesses of these tools and potential areas for improvement. We randomly selected 100 failed instances from the best tool, FastLog, and the first two authors cross-checked the causes. The analysis results are presented in Table 5. The most common failure, occurring in 56 instances, is due to Using the wrong Logging Name (i.e., Logger, LOG), indicating that incorrect or non-existent log functions are being invoked. Using Undefined Methods accounts for 21 failures, followed by Using Undefined Variables with 15 failures. Less frequent issues include Incompatible Types (3), and Unreachable Statements (1). Others mean generating the wrong syntax code, which we will not analyze.
The majority of failures (totaling 92 instances) involve undefined references to methods (21 instances), variables (15 instances), or logging names (56 instances). These undefined reference failures are primarily due to the limited context provided by existing function-level logging tools, lacking critical details on valid variables, methods, libraries, and packages relevant to the target function. Current tools are designed for function-level input, highlighting the need for logging tools to integrate better context awareness and validation checks to ensure compatibility with the existing codebase. The less frequent errors, such as incompatible types and unreachable statements, also indicate challenges in generating log statements that integrate well with the surrounding code logic. The issue of Unreachable Statements is notable because it points to a fundamental weakness in the current logging tools—specifically, their lack of understanding of the code’s control flow.
Answer to RQ2. FastLog achieved the best performance in generating compilable log statements, yet over 20% of the generated log statements still failed to be compiled. According to our analysis of compilation failure reasons, these failures primarily stem from a lack of critical contexts corresponding to the target function, e.g., valid variables, methods, libraries, packages, execution paths, and type information. To improve the reliability of automatic logging tools, it is crucial that they incorporate mechanisms to gather and utilize this additional context during the log statement generation process.
4.4. RQ3: How well do the generated log statements print logs in dynamic evaluation?
To answer RQ3, we used the dynamic evaluation dataset including code snippets paired with corresponding unit tests to generate logs. For each instance, we replaced the existing log statements in the source code with the predicted log statements from the automatic logging tools. The unit tests were then executed to observe the logs generated by the modified code.
It is important to highlight that, aside from $LEONID_{M}$, all other tools operate under the strong assumption that the given code snippet requires exactly one log statement. To highlight this inappropriate assumption, we only allow the tools one chance to predict the log statement, even in cases where multiple log statements might be needed. This experimental setup closely mirrors real-world conditions and allows for a more thorough evaluation of the tool’s ability to handle dynamic and varied logging requirements, moving beyond the oversimplified assumption that each snippet requires only one log statement. ultimately limiting tool performance in practical applications.
Table 6. The semantic similarity between logs printed by the predicted log statements and the ground truth.
Method Cosine Similarity BLEU BLEU-1 BLEU-4 ROUGE-1 ROUGE-L
$FastLog$ 0.213 16.172 24.829 15.963 24.427 23.841
$UniLog_{{cl}}$ 0.174 13.625 19.015 13.423 19.167 18.749
$UniLog_{{ds}}$ 0.130 11.607 14.674 11.556 14.002 13.869
$LANCE$ 0.099 8.201 11.342 7.971 11.235 11.033
$LEONID_{M}$ 0.072 5.466 7.946 5.303 8.095 7.944
$LEONID_{S}$ 0.044 3.220 4.908 2.955 5.099 5.012
Table 6 compares the semantic similarity between logs from source and predicted log statements. The results indicate that logs from the predicted statements significantly deviate from the ground truth, with consistently low scores across metrics like Cosine Similarity, BLEU, and ROUGE. Specifically, low similarity scores indicate that generated logs frequently fall short of matching expected outputs. Limiting each code snippet to a single log statement often sharply compromises log similarity, especially when multiple statements are needed. This limitation is particularly clear when comparing $LEONID_{S}$ and $LEONID_{M}$. Although $LEONID$ overall performs poorly, $LEONID_{M}$ stands out as the only tool capable of generating multiple log statements, which enables it to outperform $LEONID_{S}$ in similarity scores. This difference underscores the importance of tools being able to determine the appropriate number of log statements for accurate and effective logging, rather than assuming a single statement suffices.
Table 7. False Positive and False Negative Logging Rates across Methods. FPLG means the predicted log statement record logs when it should not, and FNLG means the predicted log statement does not record logs when it should.
Method FPLG Rate FNLG Rate
FastLog 9.28% 18.28%
$UniLog_{{cl}}$ 6.52% 30.59%
$UniLog_{{ds}}$ 3.21% 22.88%
LANCE 5.71% 19.29%
$LEONID_{S}$ 8.15% 8.69%
$LEONID_{M}$ 7.32% 11.6%
Table 8. FPLG and FNLG Reason Analysis. For FNLG, the major reason is beyond the execution path, and for FPLG, the major reason is lower verbosity level.
Situation Reason Number
FNLG Beyond Execution Path 35
Lower Verbosity Level 24
Wrong Code Format 4
FPLG Higher Verbosity Level 30
Beyond Execution Path 3
Total - 100
To understand the low semantic similarity scores, we examined the log generation process and found many instances where predicted logs were either redundant or missed key information present in the original logs. We quantify this issue using two metrics: False Positive Log Generation (FPLG) and False Negative Log Generation (FNLG), as reported in Table 8. For example, FastLog reports a 9.2% False Positive Log Rate (FPLR), meaning logs record redundant information in 9.2% of cases, and an 18.28% False Negative Log Rate (FNLR), indicating expected information in logs were missing in 18.28% of cases. We sampled 100 examples from FastLog, the state-of-the-art (SOTA) model, and manually analyzed the reasons for mismatches with the original logs. The results are presented in Table 8. We found that the primary reasons for failures in FNLG and FPLG differ significantly. For FNLG, the most common issue was the predicted log statements being beyond the execution path (35 cases), followed by lower verbosity levels (24 cases), and a smaller number caused by wrong code format (4 cases). In contrast, for FPLG, the main problem was higher verbosity levels (30 cases), with a few instances of log statements being beyond the execution path (3 cases). Overall, verbosity mismatches and execution path discrepancies were the dominant contributors, highlighting challenges in aligning predicted logs with actual logging requirements. The factors leading to FPLG and FNLG underscore a critical issue: while static metrics offer valuable insights into the quality of generated log statements, the actual logs are shaped by numerous contextual factors. Without a thorough understanding of the execution context, it is not possible to comprehensively evaluate the quality of log statements. Even minor discrepancies can cause significant deviations between generated logs and source logs. For instance, while the predicted log statement may capture the key information required to reflect system behavior, its effectiveness can be compromised if it is not positioned along a critical execution path or if its verbosity level is mismatched. In such cases, the log statement may fail to record essential information when key events occur. This issue majorly arises from the tools’ lack of awareness of verbosity thresholds and the control graph of the code, which limits their ability to adjust verbosity and determine appropriate log positions based on the context or execution requirements. These two limitations highlight that current logging tools lack the adaptability and context-awareness needed for effective real-world application.
Answer to RQ3. The best predicted log statements by FastLog achieve only 0.231 cosine similarity with the original logs. Many predictions record redundant information, while others miss key details. The missing key information is primarily due to the prediction being placed beyond the execution path during important events, while setting a higher verbosity level in the log statements leads to redundancy. This result highlights that automatic logging tools still have significant room for improvement.
5. Threats to Vadility
  • Construct Validity: We use BLEU to assess the quality of the generated log messages. Although text similarity metrics may not fully capture the quality of the generated text (Gros et al., 2020; Roy et al., 2021), we follow previous works (Xie et al., 2023; Xu et al., 2024; He et al., 2018a; Ding et al., 2023) to use BLEU as widely-accepted quality measures for the generated log messages. We also adapt Cosine Similarity based on TF-IDF to evaluate the quality of log content. Since log content, excluding the log header, is primarily composed of natural language, cosine similarity effectively captures the semantic meaning (Salton et al., 1975).
  • Internal Validity: The second threat to validity concerns reproducing the baseline. To minimize inconsistencies, we adapted the released models from previous work (Xie et al., 2023; Mastropaolo et al., 2022, 2023) and sought guidance from the authors of the closed-source model UniLog (Xu et al., 2024) to reproduce the tool under supervision.
  • External Validity: In this work, we evaluate the tools with 10 widely recognized Java projects in different domains. The results of each tool are examined in different requirements of logging practices. The novel evaluation method we propose is not limited to Java projects, we will explore its effects in other programming languages in the future.
6. Conclusion
This paper introduces AL-Bench, designed specifically for automatic logging tools. It includes a high-quality dataset and a novel dynamic evaluation method focused on runtime logs, addressing key limitations of prior studies and bridging the gap between real-world requirements and existing evaluation frameworks. This dynamic evaluation method assesses both the compilability of predicted log statements and their effectiveness in generating runtime logs. Using AL-Bench, we evaluate popular end-to-end automatic logging tools and find that generated log statements fail to compile in 20.1% to 83.6% of cases. Even the best predictions achieve only 0.213 cosine similarity between generated and ground-truth runtime logs. These results show that automatic logging still has a long way to go.
7. DATA AVAILABILITY
All the code and data used in our study are publicly available on https://github.com/shuaijiumei/logging-benchmark-scripts.